OpenStudy (anonymous):

look at the pattern shown below what number is missin from the pattern? 20,22,26,34,__,8

OpenStudy (lgbasallote):

20 + 2 = 22 22 + 4 = 26 26 + 8 = 34 therefore the next number would be 34 + 16

OpenStudy (anonymous):

the last number i the sequence u wrote is 82, right?

OpenStudy (anonymous):

in*

OpenStudy (lgbasallote):

82 is (34 + 16 )+ 32 :D

OpenStudy (anonymous):

yes sorry its 82

OpenStudy (anonymous):

yay cool :)

OpenStudy (lgbasallote):

so what do you think will be the answer here @Sandyp524 ?

OpenStudy (anonymous):

50

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

really cool thanks u

OpenStudy (lgbasallote):

welcome ^_^